OpenStudy (anonymous):

f(x) = 5x + 2. Find f(3) a. 12 b. 13 c. 15 d. 17

OpenStudy (anonymous):

17

OpenStudy (lalaly):

f(3)=15+2=17

OpenStudy (aroub):

5(3)+2= 15+2 17

OpenStudy (anonymous):

f(x)=5x+2 f(3)=5(3)+2 f(3)=15+2 f(3)=17